OpenStudy (anonymous):

Three consecutive integers have a sum of 87 .. What are the integers ?!

undeadknight26 (undeadknight26):

The options are endless...Any answer choices?

OpenStudy (anonymous):

No...

OpenStudy (freckles):

Let n-1 be an integer. The next integer is n. The next integer after that would be n+1. so just solve the equation: (n-1)+n+(n+1)=87

undeadknight26 (undeadknight26):

What he said xD Any questions?

OpenStudy (anonymous):

I get 2n^3=86 then what?

OpenStudy (freckles):

oh mine how did you get that

undeadknight26 (undeadknight26):

Did you tell her to divide 86 into 3 first THEN do this?

OpenStudy (freckles):

how many n's do you see here n, n , n?

OpenStudy (anonymous):

3..

OpenStudy (freckles):

n+n+n=3n

OpenStudy (freckles):

-1+1=0

OpenStudy (freckles):

\[(n-1)+n+(n+1)=87 \\(n+n+n)+(-1+1)=87 \\ 3n+0=87 \\ 3n=87\]

OpenStudy (freckles):

n+n+n=3n just like 4+4+4=3(4) or 12

OpenStudy (anonymous):

So answer is 29..30..31?

OpenStudy (freckles):

3n=87 gives n=29 n-1=? n+1=?

OpenStudy (anonymous):

28,29,30

OpenStudy (freckles):

put n to be 29 these questions I just asked you are equivalent to: n-1=29-1=28 n+1=29+1=30 yes 28+29+30=87

OpenStudy (freckles):

yep we had to find the number before and after 29 since we used (n-1) and n and (n+1) to be are integer where n we found to be 29

OpenStudy (anonymous):

Thank you .. Can u stick around and help me with some other problems?

OpenStudy (freckles):

I can probably help with one more

OpenStudy (anonymous):

Write 7/3 as a mixed number

OpenStudy (anonymous):

How do I start?

OpenStudy (freckles):

how many times does 3 go into 7? or how many 3's are in 7?

OpenStudy (anonymous):

2.3

OpenStudy (freckles):

7=3+3+1 how many 3's do you see?

OpenStudy (anonymous):

2

OpenStudy (freckles):

right so far we have: \[\frac{7}{3}=2 \frac{?}{3}\]

OpenStudy (freckles):

the question mark is the remainder

OpenStudy (anonymous):

1

OpenStudy (freckles):

7=3+3+1 right that 1 right there was left over

OpenStudy (freckles):

so the question mark is 1

OpenStudy (anonymous):

Thank you!

OpenStudy (freckles):

np